OpenStudy (anonymous):

Multiply and solve for x: 2x(x – 3) = 5

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

no

OpenStudy (xapproachesinfinity):

distribute 2x over (x-3) first then bring 5 to the other side you will have a quadratic equation and you know what to do

OpenStudy (johnweldon1993):

Right, quadratic formula is one way to go...or just complete the square \[\large 2x(x - 3) = 5\] Distribute the 2x into the parenthesis \[\large 2x^2 - 6x = 5\] Divide everything by 2 \[\large x^2 - 3x = \frac{5}{2}\] Take 1/2 the coefficient of the 'x'....square it...and add that to both sides \[\large x^2 - 3x + \frac{9}{4} = \frac{5}{2} + \frac{9}{4}\] Complete thesquare on the left \[\large (x - \frac{3}{2})^2 = \frac{19}{4}\] square root both sides \[\large x - \frac{3}{2} = \pm \sqrt{\frac{19}{4}}\] Finally add \(\large \frac{3}{2}\) to both sides \[\large x = \pm {\frac{\sqrt{19}}{2}} + \frac{3}{2}\] which breaks into 2 equations \(\large x = \frac{\sqrt{19}}{2} + \frac{3}{2}\) and \(\large x = -\frac{\sqrt{19}}{2} + \frac{3}{2}\) good to know a couple ways to do it